OpenStudy (anonymous):

@Mathematics Gary bought a car for $40,000. If V = 40,000(.85)t represents the value of the car after t years, how long will it take the car to be worth less than one-fourth of its purchase price?

OpenStudy (anonymous):

set 40000(0.85)t equal to 1/4 of its original purchase price and solve for t.

OpenStudy (jamesj):

i.e., \[ 40,000 (0.85)^t = 10,000 \]

OpenStudy (jamesj):

Make sense?

OpenStudy (anonymous):

Did you want to learn the material or just get an answer?

OpenStudy (jamesj):

...because we have a bias against just giving answers, because it's bad for you and unsatisfying for us.

OpenStudy (anonymous):

Where did 10,000 come from ? and ok...

OpenStudy (jamesj):

10,000 = 1/4 of the original price of 40,000 40,000/4 = 10,000

OpenStudy (anonymous):

ooh ok

OpenStudy (anonymous):

So how would you solve given the equation james gave you?

OpenStudy (jamesj):

This equation: \[ 40,000(0.85)^t = 10,000 \]

OpenStudy (anonymous):

Um I'm not sure on how to solve it...

OpenStudy (anonymous):

I'm assuming you're learning about logarithms?

OpenStudy (anonymous):

So @JamesJ I'm multiplying 40,000 and 0.85 ??

OpenStudy (anonymous):

Try taking the log of both sides; what do you get?

OpenStudy (anonymous):

This material is new so I have no idea of what terms you're using therefore I don't understand your question.

OpenStudy (anonymous):

what is this section of the math book teaching you? Is it logarithms?

OpenStudy (anonymous):

I got to an early college program and we were given these problems to test our knowledge and ability to teach ourselves, but I decide to come to this website to receive help in what I don't know.

OpenStudy (anonymous):

OK, so step 1 is you'd need to learn about logarithms.

OpenStudy (anonymous):

ok so that is?

OpenStudy (anonymous):

Do you know about exponentiation, e.g. what's 2^5?

OpenStudy (anonymous):

32..

OpenStudy (anonymous):

so a logarithm asks, 2 to what power is 32? i.e. log(base 2) of 32 = 5.

OpenStudy (anonymous):

This is a pretty big topic - can I send you to a tutorial on it? http://webmath.amherst.edu/qcenter/logarithms/index.html

OpenStudy (anonymous):

Sure,, I'll get help from here. Thanks for helping .

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

I'm sorry to hear you think this site is "joke" just because we didn't do your work for you.

OpenStudy (anonymous):

I'm not sorry to hear that you're worried about the wrong thing because it's not that you didn't do my work, but you obviously don't know how to do it although you did waste my time . Even after you gave me the site I still appreciated it by saying thanks so dismiss yourself with that drama.

OpenStudy (anonymous):

Of course we know how to do it. You take the log of both sides. But you said you didn't know what a log was.

OpenStudy (anonymous):

And I was referring to the new question you posted saying the site was a joke

OpenStudy (anonymous):

I didn't say anything about "we" I'm addressing you and the purpose of posting the question to receive help. I can google an site all by myself and I know exactly what you were referring to but guess what... it's my opinion and if I want to state it I will.

OpenStudy (anonymous):

Well, sure, I never said you weren't allowed to state an opinion, just said I disagreed. Do you want more help solving this?

OpenStudy (anonymous):

I know how to solve and and I'll help you if you want, but I'm not a guy who just gives out answers.